# References.bib

### v1

@@ Abstract
With the advent of accessible interfaces for interacting with large language models, there has been an associated explosion in both their commercial and academic interest. Consequently, there has also been an sudden burst of novel attacks associated with large language models, jeopardizing user data on a massive scale. Situated at a comparable crossroads in its development, and equally prolific to LLMs in its rampant growth, blockchain has emerged in recent years as a disruptive technology with the potential to redefine how we approach data handling. In particular, and due to its strong guarantees about data immutability and irrefutability as well as inherent data provenance assurances, blockchain has attracted significant attention as a means to better defend against the array of attacks affecting LLMs and further improve the quality of their responses. In this survey, we holistically evaluate current research on how blockchains are being used to help protect against LLM vulnerabilities, as well as analyze how they may further be used in novel applications. To better serve these ends, we introduce a taxonomy of blockchain for large language models (BC4LLM) and also develop various definitions to precisely capture the nature of different bodies of research in these areas. Moreover, throughout the paper, we present frameworks to contextualize broader research efforts, and in order to motivate the field further, we identify future research goals as well as challenges present in the blockchain for large language model (BC4LLM) space.
@@ 1 Introduction
From disparate areas such as software development, the solicitation of political advice, and assistance in creative writing tasks, the introduction of large language models into everyday life has occurred at unprecedented pace and scale [teubner_welcome_2023]. Consequently, many of the vulnerabilities that large language models contend with are well known and understood in the current literature, such as prompt injection [mbula_mboma_assessing_2023][shen_large_2023], hallucinations [kasneci_chatgpt_2023][zhang_sirens_2023][pan_privacy_2020][bender_dangers_2021], and data poisoning [gong_dynamic_2023][keshk_privacy-preserving-framework-based_2020]. Despite this, relatively little in terms of mitigation has been introduced to combat such weak points [yao_survey_2024][abdali_securing_2024]. The ramifications of this fact are far-reaching in regards to user experience and mounting data integrity concerns surrounding AI systems at large [winograd_loose-lipped_2022][harrer_attention_2023][lukas_analyzing_2023].
@@
For example, and representative of the variety of threats that LLMs face, is the concerning fact that LLMs are easily coaxed into revealing personally identifiable information (PII), as in the case when an LLM was persuaded into divulging undisclosed names, physical addresses, emails, phone numbers, and twitter handles associated with specific individuals to an unauthorized user [carlini_extracting_2023]. Similarly, attacks such as prompt injection can be compounded with a model’s tendency to divulge information, resulting in massive data leakage [winograd_loose-lipped_2022]. Typically, defensive responses to these threats manifest themselves in the application of established machine learning techniques, such as differential privacy strategies applied to entire corpus’ to improve privacy guarantees [abadi_deep-learning-dp][yan_protecting_2024]. While this is an important application, it is often the case that DP guarantees are not enough to fully ensure data privacy in a large language model [lukas_analyzing_2023] due to DP’s ability to protect primarily ”by whom” data is contributed, rather than ”about whom” the data is focused on. Additionally, another common attempt to tackle the data privacy problem in LLMs is the use of federated learning (FL) in the training process. This technique distributes training across multiple nodes to create a decentralized environment for model data to exist within [mcmahan_communication-efficient_2017]. Naturally, this lends itself to further obscuring sensitive information in a model’s corpus. However, it has been shown that by taking model weights or gradients, original data from the model can still be reconstructed [liu_federated_2021]. Additionally, federated learning solutions are susceptible to many of the same types of attacks as large language models, such as single-point-of-failure attacks or man-in-the-middle attacks [qu_blockchain-enabled_2023]. This trend of typical machine learning solutions failing to exhaustively defend against the range of attacks now affecting LLMs continues across multiple traditional threat/defense models [shayegani_survey_2023, zhang_defending_2024, yan_protecting_2024].
It is clear that to counter new threats arising from large language models there is a simultaneous need for new technologies to be introduced into the space. Situated in a similar position to LLMs in their rapid emergence and adoption are blockchain systems. They have the ability to ensure data integrity via various tamper-evident mechanisms, introduce a high degree of confidentiality into otherwise centralized systems, and allow for data provenance guarantees through traceable and auditable data [deshpande_distributed_nodate][cao_blockchain_2023][mingxiao_review_2017][ali_comparative_2021]. These qualities of blockchain systems place themselves at an ideal juncture to be used for bolstering the robustness of large language models. This integration creates the necessary conditions to allow for stronger privacy preservation, enhanced inference checks, anti-adversarial attack technologies, and similar defenses to be incorporated into the design of large language models.
This unification of technologies has become a hot topic of research within both blockchain and large language model oriented communities over the last several years. As a field in its infancy, blockchain for large language models (BC4LLM) demands rigorous analysis so that it may progress unhindered from any limitations that may arise from the combination of two newly introduced fields, especially given their relative complexity and various nuances. Towards this end of progress in the fields of large language models and blockchain systems, we present this survey paper as a vehicle for researchers to better grasp the state of the two fields and especially understand the combination of the two technologies in the vein of how blockchain can better serve large language model systems. To clarify our goals with this paper, we introduce four research questions that motivate the entirety of the paper and generalize our aim in writing this survey. They are as follows:
@@
Of note is our focus specifically on how blockchain systems may impact large language model security and safety. We limit the scope of this survey to these terms in order to provide finer-grained analysis and categorization of seemingly disparate works - and to more clearly spur research advances in specific directions. Indeed, we typically see attacks on LLMs occur via malicious third parties which exploit system vulnerabilities (security) [anderson_is_2024][xue_badrag_2024][zhang_defending_2024][hu_membership_2022][liu_prompt_2024] or as passive issues embedded in the structure of a LLM which places users at risk with no malicious outside influence (safety) [fraga-lamas_fake_2020][wan_kelly_2023][seneviratne_blockchain_2022][yazdinejad_making_2020]. It is because of this encompassing distinction that we offer our analysis of the blockchain for LLM (BC4LLM) space in the context of security and safety. Moreover, we further bolster the significance of this distinction with definitions of security and safety in the context of large language models. To the best of our knowledge, we are the first paper to rigorously introduce such definitions as we aim to lay a foundation upon which future BC4LLM works can build. It is also worth pointing out our contributions towards further defining privacy measures in the form of active privacy and passive privacy efforts, modeled after Yan et al’s survey on bolstering data privacy [yan_protecting_2024].
Table 1: Overview of Existing Related Surveys. Reviews concerning blockchain and large language models take many different forms and facilitate different functions within the broader context of BC4LLMs literature. However, a holistic review of BC4LLMs is absent from the current state of the field. For example, while many survey papers have focused to some degree on general threats affecting LLMs, fewer have considered LLM-specific threats or approached these problems from the direction of using blockchain technology. In this table, we document to what degree several important components of our survey are present in different reviews. In particular, we examine if the background is given on relevant subjects, if a model of threat categorization is introduced, if definitions of security and safety are proposed, whether security and/or safety in regards to BC4LLMs is explored, if future BC4LLMs work is probed, and whether or not the survey instead focuses on LLM for Blockchain (LLM4BC) instead of BC4LLM.
  denotes a full discussion of the related topic,   a partial discussion, and   indicates the topic was absent from the review.
@@
Source LLM and BC Background Threat Model Definitions Security in BC4LLMs Safety in BC4LLMs Future Work LLM for Blockchain
@@
To separate ourselves, and our analysis of BC4LLM through the lens of safety and security, from other similar works, we look at several other closely related reviews. He et al. [he_large_2024] examine the relationship between large language models and blockchain in analyzing how large language models can further enhance blockchain systems in the LLM4BC space. Mboma et al. [mbula_mboma_assessing_2023] provide an exploratory review of general integrations between blockchain and large language models, which is similar to Heston’s analysis of integrating the two technologies in the sphere of telemedicine [f_heston_perspective_2024]. In the case of Salah, et al. [salah_blockchain_2019], Bhumichai et al. [bhumichai_convergence_2024], and Dinh et al. [dinh_ai_2018], we see the exposition of potential and existing technologies between blockchain and artificial intelligence in general. In summary, existing reviews that contend specifically with the blockchain and large language model space are not focused on the direction in which they apply the technologies. On the other hand, reviews that concern blockchain and AI as a whole lose the benefits of tighter granularity and focus. We present an overview of related works in table 1, which juxtaposes the above papers’ contents with our specific focuses, stressing the differentiating factors that we introduce into the space.
1.1 Contribution
We outline the exact contributions of this survey paper, and highlight their impact in answering our research questions. They are as follows:
@@ 1.1 Contribution
    We present several key contributions in the form of various frameworks, definitions, and compiled resources. Most prominently, we introduce a taxonomy of blockchain for large language Models in Figure 3. This taxonomy aims to succinctly explain the relevant interactions between various blockchain components and corresponding large language model vulnerabilities [RQ1][RQ3]. To further motivate and contextualize this taxonomy as well as our general discussion of existing literature, we introduce two definitions of safety and security as they apply to large language models [RQ2]. Moreover, we also provide multiple datasets relevant to BC4LLM in order to provide future researchers in the area with the tools to expand upon connections highlighted by the taxonomy and motivated by our definitions [RQ4].
@@
    We also highlight several other important components of our paper, intended to support our main contributions but still relevant in their own regard as pertinent artifacts of the BC4LLMs space. One such artifact is our definitions of specific areas that are found within our definition of safety, further expanded upon in Table 3 [RQ2][R3]. These definitions add further weight to our definition of safety for large language models. We go on to bolster our definitions of both safety and security by purposefully focusing on the issue of privacy in the context of security, consequently reaffirming two terms as introduced by Yan et al. [yan_protecting_2024]: passive and active privacy [RQ1][RQ3]. Also relevant in their own right but primarily proposed as supporting contributions are our contextualization of LLMs as a sub-field of various AI areas [RQ1], and our succinct taxonomic overview of blockchain components [RQ1][RQ3].
@@
    Finally, we make note of the overarching goal of our paper, which also stands as the most critical contribution we offer. That is, we present this paper in part as a holistic literature review in Section 4. In it, we categorize relevant fields to present novel insights into the space, classify research efforts in several interrelated respects, relate all BC4LLM research projects to LLM safety and security, and as a result capture an accurate and informative picture of how blockchain can benefit large language models [RQ1][RQ2][RQ3][RQ4].
The remainder of the survey is organized as follows: Section 2 introduces the background of blockchain technology and large language models. Section 3 consists of our methodology; we analyze how we filtered works for this review, and introduce the relevant definitions which motivate our analysis of the current literature. We also share our model of threat categorization, similar to Yao et al.’s categorization [yao_survey_2024]. In Section 4, we present the comprehensive literature review of the BC4LLM space in the context of safety and security. We then scrutinize works within these areas in relation to our BC4LLM taxonomy. Within Section 5, we present relevant datasets to the BC4LLM space in order to supply future researchers with both well-known and more obscure datasets. Afterwards, in Section 6, we discuss important challenges that exist between blockchain technology and large language models which fundamentally hinder progress in the field at large. Section 7 proposes future research directions in the BC4LLM space, and section 8 concludes the paper, which both summarizes our efforts in surveying the BC4LLM area and provides further insight into the state of BC4LLM as a whole.
@@ 2.1 Blockchain
Since the introduction of Bitcoin as a decentralized currency by Satoshi Nakamoto in 2008 [nakamoto_bitcoin_2008] there has been a subsequent explosion of academic and commercial interest in its underlying blockchain technology. Additionally, and as highlighted by the introduction of Vitalik Buterin’s Ethereum blockchain in 2014 [buterin_ethereum_2014], there has been a particular focus on blockchain’s potential applications in fields entirely disparate from digital currencies. The interest in blockchain, or distributed ledger technologies, stems from its guarantees about data sovereignty, transparency, and relative permanence. Concisely, these properties are often referred to as immutability and irrefutability. Ranging from many diverse fields such as health care record management, digital identity management, or tax auditing, these properties are widely applicable and highly desirable, even though the mechanisms through which we achieve them can be somewhat complex and opaque. In light of the oftentimes convoluted nature of blockchain systems, we introduce blockchain to the reader in a piecemeal fashion in order to emphasize the modular, yet interconnected nature of such systems. Figure 1 represents an overview of our characterization of blockchain systems in general. We purposefully exclude certain components such as the incentive mechanism, or wallets, as they are beyond the scope of our analysis of blockchain as a means to serve large language models.
@@ 2.1.1 Blockchain Components
Consensus Protocol. Of particular interest to the BC4LLM space, and arguably the most fundamental component within a blockchain, the consensus protocol is the governing system that controls how data is added to a blockchain’s ledger. At its core is the consensus mechanism, which both ensures the validity of proposed data and fosters an environment of accountability, so that nodes submitting invalid information may be penalized accordingly. For example, the Proof of Work (PoW) consensus mechanism [nakamoto_bitcoin_2008] is by far the most widely known. In it, nodes must solve a complex mathematical equation in order to gain rights to propose data for the blockchain. When such a node submits new data, it is scrutinized by every other node in the system. If the data is malicious, or untruthful, the proposal is rejected and the corresponding processing power performed by the malicious node has effectively been wasted, as that node will not receive the incentive, a Bitcoin reward. The underlying ideas of accountability, certain nodes being selected as ’block proposers’, and the ’proof’ of the ability to submit information to the chain are central ideas in consensus protocols across blockchains with different consensus protocols [nguyen_proof-of-stake_2019].
@@
Verifiable Ledger. At a blockchain’s core sits the verifiable ledger, a repository of data bolstered by a secure way of maintaining the integrity of that data. Of note is the particular technique through which data itself is verified on the ledger: the Merkle tree [merkle_digital_1988], or a variation thereof. Typically implemented as a ground-up binary tree, data is stored in leaf nodes, with hashed pointers of that data cascading up the tree. This structure results in a comprehensive ’Merkle root’, a hash pointer consisting of all the other hash pointers in lower levels of the tree, which is ultimately based on the data stored in the leaf nodes. This technique ensures the integrity of information in the leaf nodes, as any alteration to the data is instantly reflected in the Merkle root. Likewise, new additions to the Merkle tree can be checked against previous states of the tree via a recalculation of the Merkle root accounting for the new transactions. This technique, complementing the verifiable ledger, is often the key to LLM data provenance and traceability solutions that rely on blockchain technology.
@@
Decentralized Network. Critically, blockchain’s are decentralized networks. That is, no central server or group of servers may assume control of the network in a way that would compromise the network’s state of trustlessness. This is achieved through multiple avenues, such as the aforementioned consensus protocol, the distribution of the verifiable ledger among a large number of independent nodes, and the accessibility of a given blockchain’s network. [deshpande_distributed_nodate] In this way, no users in the network are required to trust any other user. This fundamental aspect of blockchain is responsible for already realized and potential advancements with LLMs concerning areas such as RAG, the training process, and even supply chain issues.
@@
Layer 2 Technologies. Apart from the fundamental components found within all blockchains themselves, there exist several external architectures that interface with blockchains and further enhance their applicability. Typically, these are referred to as layer 2 technologies, as they sit a ‘layer’ above the ‘layer 1’ blockchain. Increasingly relevant as blockchain’s influence grows, layer 2 solutions are a burgeoning area with numerous novel research directions. Most prominent among these is the space concerning smart contracts, scripts that rely on a blockchain’s security guarantees in order to facilitate off-chain transactions [zheng_overview_2020]. Also of note in the layer 2 field, and sometimes combined with the efficacy of smart contracts, are zero-knowledge rollups. Often used to strengthen scalability, zero-knowledge rollups batch unproposed transactions together, and instead of submitting the transactions themselves, submit proof that the transactions are indeed valid [sun_survey_2021]. This allows for transactions to be added on-chain without the need for every full node to redo the calculations found within those transactions. This area of layer 2 technologies is pivotal as it relates to BC4LLM - layer 2 has the necessary dynamism to react quickly to new and emerging LLM vulnerabilities.

### v2

@@ Abstract
With the growing development and deployment of large language models (LLMs) in both industrial and academic fields, their security and safety concerns have become increasingly critical. However, recent studies indicate that LLMs face numerous vulnerabilities, including data poisoning, prompt injections, and unauthorized data exposure, which conventional methods have struggled to address fully. In parallel, blockchain technology, known for its data immutability and decentralized structure, offers a promising foundation for safeguarding LLMs. In this survey, we aim to comprehensively assess how to leverage blockchain technology to enhance LLMs’ security and safety. Besides, we propose a new taxonomy of blockchain for large language models (BC4LLMs) to systematically categorize related works in this emerging field. Our analysis includes novel frameworks and definitions to delineate security and safety in the context of BC4LLMs, highlighting potential research directions and challenges at this intersection. Through this study, we aim to stimulate targeted advancements in blockchain-integrated LLM security.
@@ 1 Introduction
The widespread application of large language models (LLMs) has progressed at an unprecedented pace and scale in our daily lives [teubner_welcome_2023]. Such a widespread application exposes several vulnerabilities inherent to LLMs, such as data poisoning [keshk_privacy-preserving-framework-based_2020, gong_dynamic_2023], prompt injections [mbula_mboma_assessing_2023, shen_large_2023], and hallucinations [kasneci_chatgpt_2023, zhang_sirens_2023, pan_privacy_2020, bender_dangers_2021]. For example, prompt injections can exploit a model’s propensity to disclose information, resulting in significant data leakage, like leaking personally identifiable information (PII) to an unauthorized user [winograd_loose-lipped_2022, carlini_extracting_2023]. Although numerous studies have attempted to address these issues [yao_survey_2024, abdali_securing_2024], there remains no effective mitigation strategies capable of addressing growing concerns about these issues in LLMs [winograd_loose-lipped_2022, harrer_attention_2023, lukas_analyzing_2023]. Typically, defensive strategies against these threats are implemented through established machine-learning methods, such as applying differential privacy (DP) techniques to the entire dataset to enhance privacy protections [abadi_deep-learning-dp, yan_protecting_2024]. While DP strategies are one of the crucial applications, these strategies cannot fully guarantee data privacy in LLMs [lukas_analyzing_2023] due to DP’s ability to protect primarily “by whom” data is contributed, rather than “about whom” the data is focused on. Additionally, another common attempt to tackle the data privacy problem in LLMs is to apply federated learning (FL) techniques in the training process by distributing model training across multiple nodes to create a decentralized environment [mcmahan_communication-efficient_2017]. Naturally, this lends itself to further obscuring sensitive information in a model’s corpus. However, it has been shown that by taking model weights or gradients, original data from the model can still be reconstructed [liu_federated_2021]. What’s worse, federated-learning approaches are susceptible to many of the same types of attacks as large language models, such as single-point-of-failure attacks or man-in-the-middle attacks [qu_blockchain-enabled_2023]. This trend of typical approaches failing to exhaustively defend against the range of attacks now affecting LLMs continues across multiple traditional threat/defense models [shayegani_survey_2023, zhang_defending_2024, yan_protecting_2024].
To address the limitations of the above-mentioned methods and further enhance data privacy, blockchain technology has emerged as a promising solution [mingxiao_review_2017]. It ensures data integrity through various tamper-evident mechanisms, introduces a high level of confidentiality to otherwise centralized systems, and guarantees data provenance by enabling traceable and auditable records [deshpande_distributed_nodate, cao_blockchain_2023, ali_comparative_2021]. These benefits can significantly strengthen the robustness of large language models. Integrating blockchain technology lays the foundation for stronger privacy protection, enhanced inference validation, defenses against adversarial attacks, and other security measures to be incorporated into the design of large language models. This overlapping research direction is still in its early stages. To facilitate a deeper understanding of the current landscape for emerging researchers, we conduct a comprehensive literature review in this paper to explore how blockchain technology can better serve large language models (BC4LLMs). Overall, our objectives in this survey paper are to address the following four research questions:
@@
Notably, we focus specifically on how blockchain systems may impact large language models’ security and safety. By narrowing our scope to these dimensions, we aim to provide a more detailed analysis and categorization of seemingly disparate works, thereby encouraging targeted research advances in specific directions. In general, attacks on LLMs typically manifest in two primary ways: as direct exploitations by malicious third parties that capitalize on system vulnerabilities (security) [anderson_is_2024, xue_badrag_2024, zhang_defending_2024, hu_membership_2022, liu_prompt_2024] and as inherent risks embedded within LLM structures that expose users to potential harm without external malicious influence (safety) [fraga-lamas_fake_2020, wan_kelly_2023, seneviratne_blockchain_2022, yazdinejad_making_2020]. We base our analysis of the blockchain for LLMs (BC4LLMs) on this critical distinction between security and safety, a distinction that we underscore through explicit definitions contextualized for LLMs. To the best of our knowledge, this is the first study to rigorously define these terms in the BC4LLMs context, providing a foundation for subsequent work in this area. Furthermore, we contribute to the discourse on privacy in LLMs by delineating active and passive privacy efforts, modeled after a survey about data privacy [yan_protecting_2024].
To distinguish our analysis of BC4LLMs from other similar works through the lenses of safety and security, we compare several reviews in this domain. He et al. [he_large_2024] examine the relationship between LLMs and blockchain in analyzing how LLMs can further enhance blockchain systems. Mboma et al. [mbula_mboma_assessing_2023] provide an exploratory review of general integrations between blockchain and large language models, which is similar to Heston’s analysis of integrating the two technologies in telemedicine [f_heston_perspective_2024]. Additionally, Salah et al. [salah_blockchain_2019], Bhumichai et al. [bhumichai_convergence_2024], and Dinh et al. [dinh_ai_2018] provide an overview of potential and existing technologies between blockchain and artificial intelligence in general. In short, current reviews that specifically address blockchain and LLMs lack a clear focus on the specific applications of these technologies, whereas broader reviews that encompass blockchain and AI sacrifice the depth of analysis. To close this gap, we present an overview of related surveys in Table 1, which juxtaposes the above papers’ contents with our specific focuses, highlighting the distinction in our study.
Table 1: Overview of Existing Related Surveys. We compare related surveys about blockchain techniques and LLMs from various perspectives, such as background, threat model, definition, security, safety, etc. In particular, we are interested in investigating whether (i) relevant subjects are discussed in the background section, (ii) a model of threat categorization is introduced, (iii) definitions of security and safety are proposed, (iv) security and/or safety with regards to BC4LLMs is explored, (v) future BC4LLMs work is probed, (vi) the survey focuses on LLMs for Blockchain. We denote , , and  as a full, partial, and no discussion of the corresponding items.
@@
Source LLMs and BC Background Threat Model Definitions Security in BC4LLMs Safety in BC4LLMs Future Work LLMs for Blockchain
@@
We outline our main contributions and highlight the impact to answer our research questions as follows:
@@
    In this work, we first contribute a series of frameworks, definitions, and compiled resources. Most prominently, we propose a new taxonomy about applying blockchain techniques for LLMs in Figure 3. Through the proposed taxonomy, we aim to succinctly explain the relevant interactions between the blockchain techniques and corresponding LLMs’ vulnerabilities [RQ1][RQ3]. To further contextualize this taxonomy and ground our discussion of existing literature, we propose two foundational definitions of safety and security specific to LLMs [RQ2]. Moreover, we provide a collection of datasets relevant to BC4LLMs, equipping future researchers with resources to build on the connections delineated by our taxonomy and informed by our definitions [RQ4].
@@
    We also highlight additional components of our paper that, while supporting our main contributions, serve as valuable artifacts in the BC4LLMs space. One such artifact is our definition of specific areas within the broader concept of safety, further detailed in Table 3 [RQ2][R3]. These definitions reinforce our conceptualization of safety for LLMs. To enhance our definitions of both safety and security, we specifically address privacy within the security context, reaffirming two terms introduced by Yan et al. [yan_protecting_2024]: passive and active privacy [RQ1][RQ3]. Besides, our contextualization of LLMs within various AI sub-fields [RQ1] and our concise taxonomic overview of blockchain components [RQ1][RQ3] hold intrinsic value as distinct contributions to the field.
@@
    Last, we conduct a comprehensive literature review in Section 4, where we classify research works across several interrelated domains, offer novel insights within these categorized domains, and align all BC4LLMs research projects with LLMs’ safety and security. By conducting this review, we provide an informative perspective on the potential of utilizing blockchain techniques to enhance LLMs [RQ1][RQ2][RQ3][RQ4].
The remaining sections are organized as follows: In Section 2, we introduce the background of blockchain technology and large language models. In Section 3, we describe our methodology, including the criteria used to filter works for this review and the relevant definitions that guide our analysis of the current literature. We also share our model of threat categorization, which aligns with the categorization proposed by Yao et al. [yao_survey_2024]. In Section 4, we conduct a comprehensive literature review of BC4LLMs in safety and security, examining key works in relation to our proposed taxonomy. In Section 5, we present datasets relevant to BC4LLMs. In Section 6, we address key challenges at the intersection of blockchain technology and LLMs that hinder advancement in this area. In Section 7, we discuss future research directions within the field of BC4LLMs. Finally, in Section 8, we summarize our efforts, providing a holistic view of the current progress of BC4LLMs.
@@ 2.1 Blockchain
Since Satoshi Nakamoto [nakamoto_bitcoin_2008] introduced Bitcoin as a decentralized currency in 2008, there has been a subsequent explosion of academic and commercial interest in its underlying blockchain technology [liang2021omnilytics, song2024unveiling, song2024advancing]. Additionally, and as highlighted by the introduction of Vitalik Buterin’s Ethereum blockchain in 2014 [buterin_ethereum_2014], there has been a particular focus on blockchain’s potential applications in fields entirely disparate from digital currencies. The interest in blockchain, or distributed ledger technologies, stems from its guarantees of data sovereignty, transparency, and relative permanence. Concisely, these properties are often referred to as immutability and irrefutability. Ranging from many diverse fields such as health care record management, digital identity management, or tax auditing, these properties are widely applicable and highly desirable, even though the mechanisms through which we achieve them can be somewhat complex and opaque. In light of the oftentimes convoluted nature of blockchain systems, we introduce blockchain to the reader in a piecemeal fashion in order to emphasize the modular, yet interconnected nature of such systems. Figure 1 represents an overview of our characterization of blockchain systems in general. We purposefully exclude certain components such as the incentive mechanism, or wallets, as they are beyond the scope of our analysis of blockchain as a means to serve large language models.
@@ 2.1.1 Blockchain Components
Consensus Protocol. Of particular interest to BC4LLMs, and arguably the most fundamental component within a blockchain, the consensus protocol is the governing system that controls how data is added to a blockchain’s ledger. At its core is the consensus mechanism, which both ensures the validity of proposed data and fosters an environment of accountability, so that nodes submitting invalid information may be penalized accordingly. For example, the Proof of Work (PoW) consensus mechanism [nakamoto_bitcoin_2008] is by far the most widely known. In it, nodes must solve a complex mathematical equation in order to gain rights to propose data for the blockchain. When such a node submits new data, it is scrutinized by every other node in the system. If the data is malicious, or untruthful, the proposal is rejected and the corresponding processing power performed by the malicious node has effectively been wasted, as that node will not receive the incentive, a Bitcoin reward. The underlying ideas of accountability, certain nodes being selected as ’block proposers’, and the ’proof’ of the ability to submit information to the chain are central ideas in consensus protocols across blockchains with different consensus protocols [nguyen_proof-of-stake_2019].
@@
Verifiable Ledger. At a blockchain’s core sits the verifiable ledger, a repository of data bolstered by a secure way of maintaining the integrity of that data. Of note is the particular technique through which data itself is verified on the ledger: the Merkle tree [merkle_digital_1988], or a variation thereof. Typically implemented as a ground-up binary tree, data is stored in leaf nodes, with hashed pointers of that data cascading up the tree. This structure results in a comprehensive ‘Merkle root’, a hash pointer consisting of all the other hash pointers in lower levels of the tree, which is ultimately based on the data stored in the leaf nodes. This technique ensures the integrity of information in the leaf nodes, as any alteration to the data is instantly reflected in the Merkle root. Likewise, new additions to the Merkle tree can be checked against previous states of the tree via a recalculation of the Merkle root accounting for the new transactions. This technique, complementing the verifiable ledger, is often the key to LLM data provenance and traceability solutions that rely on blockchain technology.
@@
Decentralized Network. Critically, blockchains are decentralized networks. That is, no central server or group of servers may assume control of the network in a way that would compromise the network’s state of trustlessness. This is achieved through multiple avenues, such as the aforementioned consensus protocol, the distribution of the verifiable ledger among a large number of independent nodes, and the accessibility of a given blockchain’s network. [deshpande_distributed_nodate] In this way, no users in the network are required to trust any other user. This fundamental aspect of blockchain is responsible for already realized and potential advancements with LLMs concerning areas such as RAG, the training process, and even supply chain issues.
@@
Layer 2 Technologies. Apart from the fundamental components found within all blockchains, several external architectures interface with blockchains and further enhance their applicability. Typically, these external architectures are referred to as layer 2 technologies, as they sit a ‘layer’ above the ‘layer 1’ blockchain. Increasingly relevant as blockchain’s influence grows, layer 2 solutions are a burgeoning area with numerous novel research directions. Most prominent among these are smart contracts, scripts that rely on a blockchain’s security guarantees to facilitate off-chain transactions [zheng_overview_2020]. Also, in layer 2, zero-knowledge rollups are often combined with the efficacy of smart contracts. Often used to strengthen scalability, zero-knowledge rollups batch unproposed transactions together, and instead of submitting the transactions themselves, submit proof that the transactions are indeed valid [sun_survey_2021]. This allows for transactions to be added on-chain without the need for every full node to redo the calculations found within those transactions. This area of layer 2 technologies is pivotal as it relates to BC4LLMs - layer 2 has the necessary dynamism to react quickly to new and emerging LLM vulnerabilities.